# SN 2024aecx: A double-peaked rapidly evolving Type IIb supernova at 11 Mpc

### v1

@@ Abstract
We present the results of low-resolution spectroscopic and densely sampled multiband simultaneous optical imaging (ugi and vrz bands) follow-up of supernova (SN) 2024aecx. The photometric data is supplemented with Swift/UVOT and ATLAS survey observations. The SN was discovered in the spiral galaxy NGC 3521 (distance 11 Mpc) within a day after the explosion. The early spectra of SN 2024aecx show a weak signature of hydrogen lines, which disappeared in 30 days after the explosion. Light curves in all bands show a distinct feature of two peaks, and the first peak is likely due to the shock cooling emission. The early phase light curve evolution of SN 2024aecx has similarity with the typical Type IIb events, but the decay rate in different bands (e.g., = 1.60 0.05 mag, -band) is significantly faster in the post-peak phase. It attained the secondary maximum in 19 days (-band) with a peak absolute magnitude of Mg = –17.94 0.10 mag. The color evolution of SN 2024aecx is displaying a red-blue-red trend between days 8 to 40. The analytical model fitting to the light curves reveals an envelope mass and progenitor radii in the range of 0.03–0.24 and 169–200 , respectively. Modeling of the pseudo-bolometric light curve suggests that synthesized 56Ni in the explosion was 0.15 M☉ with ejecta mass and kinetic energy of 0.7 M☉ and 0.16 1051 erg, respectively. The observational properties and modeling indicate that the SN 2024aecx progenitor belongs to the extended progenitor category.
@@ 1 Introduction
The evolution of massive stars (8 M⊙) ends when their nuclear fuel is exhausted after the formation of iron group elements in the core (Heger et al., 2003; Woosley & Weaver, 1986). At this stage, the stellar core collapses (Woosley & Janka, 2005) due to the strong gravitational force on a very short time scale (of the order of a second), and eventually the star explodes as a core-collapse supernova (CC-SN). Various stellar parameters (e.g., mass, metallicity, mass loss, rotation, binary interaction, etc.) may significantly affect the observational features such as light curves and spectra. Accordingly, the observational properties displayed by these transients at early or maximum light phases make a decisive basis for their classification and subclassification (Filippenko, 1997). A subgroup (Type IIb, Ib, Ic, and Ic-BL) of these events is termed as ‘stripped-envelope supernovae, SE-SNe’ (Minkowski, 1941; Clocchiatti & Wheeler, 1997; Gal-Yam, 2017; Modjaz et al., 2019).
@@
In SE-SNe, the hydrogen and/or helium envelopes of their progenitor are partially or completely stripped off before the explosion. In particular, hydrogen lines do not appear in SNe Ib, but helium lines are dominant. SNe Ic exhibit neither of these lines in their spectra, and SNe Ic-BL (broad line) are characterized with broad line velocities (15000 km s-1). Type IIb events demonstrate intriguing properties in terms of both spectroscopic and light curve evolution. During the early phases, the hydrogen lines are clearly visible, which disappear progressively in a few weeks (Filippenko, 1988), and the helium lines dominate the spectra at the later phases (similar to SNe Ib). In addition, double peaks have been observed in the light curves of a handful of IIb SNe, e.g., SN 1993J (Richmond et al., 1994; Barbon et al., 1995), SN 2011dh (Arcavi et al., 2011), SN 2011fu (Kumar et al., 2013), SN 2013df (Morales-Garoffolo et al., 2014; Van Dyk et al., 2014), SN 2016gkg (Tartaglia et al., 2017a; Bersten et al., 2018a), SN 2020bio (Pellegrino et al., 2023), SN 2022hnt (Farah et al., 2025), SN 2024uwq (Subrayan et al., 2025). While the initial peak is mainly interpreted by the cooling of the extended hydrogen-rich envelope in the post-shock breakout phase (see Falk & Arnett, 1977a; Bersten et al., 2012a; Nakar & Piro, 2014; Piro, 2015; Sapir & Waxman, 2017; Dessart et al., 2018), the secondary peak is mainly powered due to energy released in radioactive decay of 56Ni 56Co 56Fe (Arnett, 1980). However, a few more scenarios were also proposed to explain the SNe IIb first peak: double 56Ni distribution (Orellana & Bersten, 2022), Thomson scattering and chemical mixing in the SN ejecta (Park et al., 2024). The double-peak light curve features of SNe IIb have similarities with some ‘Ca-rich’ events which show strong [Ca II] 7291, 7324 Å emission compared to [O II] 6300, 6364 Å (Filippenko et al., 2003; Kawabata et al., 2010; Perets et al., 2010; Kasliwal et al., 2012). The shape of the first peak can infer the radius and mass of the progenitor (Nakar & Piro, 2014). Chevalier & Soderberg (2010) suggested the extended and compact progenitors of SNe IIb that can produce double-peaked and single-peaked light curves, respectively.
@@
From the theoretical modeling aspect, Type IIb SNe progenitors can have both single and binary evolution channels. The radiation-driven stellar wind plays a crucial role in stripping the outer envelope (Crowther, 2007; Georgy, 2012; Yoon, 2015) in single massive star models and mass transfer via a binary companion is a preferred path in the binary stellar evolution models (Podsiadlowski, 1992; Nomoto et al., 1995; Claeys et al., 2011; Smith, 2014; Sravan et al., 2020; Dessart et al., 2024). However, the observational manifestation supports the binary progenitor, which is evident in several direct imaging (e.g., Maund et al., 2004; Folatelli et al., 2014; Ryder et al., 2018; Van Dyk et al., 2014, and references therein).
@@
Thanks to modern survey facilities, which are populating the transient phase-space with traditionally known and rare events. Consequently, large sample-based qualitative investigations have been possible in recent years (Taddia et al., 2015a; Liu et al., 2016; Lyman et al., 2016; Prentice et al., 2016; Taddia et al., 2018a; Prentice et al., 2019; Shivvers et al., 2019; Williamson et al., 2019; Ayala et al., 2025). Better cadence and sampling are equally useful to corroborate whether there is a transition between some SE-SNe, which were classically divided into different types, and recent studies point toward a continuum between them. Especially in a few cases, SNe Type IIb and Ib indicate a thin boundary (e.g., Milisavljevic et al., 2013; Gangopadhyay et al., 2023; Dong et al., 2024a; Yesmin et al., 2025). Therefore, for a better understanding of the observational properties and the progenitor, it is important to study more events in detail, having both common and/or peculiar features.
@@
NGC 3521 and SN 2024aecx
NGC 3521 is a late-type spiral galaxy SAB(rs)bc with a mixed barred and inner ring morphology (c.f. Third Reference Catalogue of Bright Galaxies, de Vaucouleurs et al., 1991). Several groups have performed multi-wavelength studies of this galaxy due to its proximity (distance = 11.40 0.56 Mpc, red-shift independent)111https://ned.ipac.caltech.edu/. It has a star formation rate of 2.1 M☉ year-1 (Leroy et al., 2008), and because of the similarity in structural parameters (stellar mass, rotation velocity, optical radius), it is considered as a near structural analogue to the MW (see Pilyugin et al., 2025, and references therein). The spectroscopic decomposition indicates that NGC 3521 is dominated by the intermediate stellar population (Coccato et al., 2018). Some of the parameters of NGC 3521 are listed in Table 1.
To examine our data pipeline, implementation of image subtraction and calibration, etc., we initiated a survey of the nearby galaxies, star clusters, and several pilot fields. Consequently, the field of NGC 3521 was observed in multi-bands at various epochs since 2023 January 17. Interestingly, the latest Mephisto observations of NGC 3521 were made on 2024 December 09, i.e., about 7 days before the SN discovery (see Section 3.1).
SN 2024aecx (ATLAS24rkq) was discovered on MJD 60660.56 (UT 2024 December 16.56) by ATLAS group in the galaxy NGC 3521 at coordinates (J2000): RA = 11h 05m 4955 DEC = – 00° 02′ 0544 (Tonry et al., 2024; Stevance et al., 2024). It is one of the very early discovered (within a day) SN after the explosion. The discovery magnitude was 14.68 0.01 (orange-ATLAS). The last non-detection was reported on MJD 60659.35 (UT 2024 December 15.35) with a magnitude of 18.56 mag (orange-ATLAS). AT 2024aecx observations (at MJD = 60660.7) from the Lulin Observatory indicated a fading nature and blue colors ( = – 0.6 mag, = – 0.3 mag) of AT 2024aecx (Chen et al., 2024a). Forced photometry with ATLAS (orange filter) confirmed the fast decline (Pérez-Fournon et al., 2024) within a few hours after the explosion. We consider the explosion epoch of SN 2024aecx as JD 2460659.95 (2024 December 15.95), the midpoint of the last non-detection, and the discovery date. The considered explosion time matches reasonably with the estimation from the analytical model fitting as discussed in Section 4. On 2024 December 17, the SN was classified as a Type IIb based on spectra taken from the SNIFS and GMOS instruments mounted on the 88-inch University of Hawaii telescope (Hinkle, 2024) and Gemini-North telescope (Andrews et al., 2024a). However, Gemini-N/GMOS spectroscopy on 2024 December 19 reclassifies it as Type Ic SN (Andrews et al., 2024b).
@@
Parameters | Value | Reference |
SN 20124aecx: | ||
RA (J2000) | 11h 05m 4955 | 1 |
DEC (J2000) | –00° 02′ 0544 | 1 |
Explosion epoch (UT) | 2024 December 15.95 | Section 1 |
(MJD 60659.95) | ||
Discovery date (UT) | 2024 December 16.56 | 1 |
(MJD 60660.56) | ||
0.45 0.10 mag | Section 3.2 | |
NGC 3521 (To Check): | ||
Type | SAB(rs)bc | 2 |
Redshift | 0.002665 0.000002 | 2 |
Distance | 11.40 0.56 Mpc | 2 |
Distance modulus | 30.28 0.11 mag | 2 |
Absolute magnitude (MB) | –21.25 0.36 | 3 |
@@ 2.1 Photometry
The optical photometric observations of SN 2024aecx were acquired with the 1.6 m Mephisto telescope, a unique wide-field multichannel observing facility. Mephisto is located at the Lijiang Observatory (IAU code: 044) of Yunnan Astronomical Observatories, Chinese Academy of Sciences (Wang et al., 2019a), and the operational management is led by the South-Western Institute for Astronomy Research, Yunnan University. Mephisto can perform simultaneous observations of a field in three optical bands at a time ( or ), which is a unique feature, especially for transient studies. The wavelength coverage of the filters , , , , , and is 320–365 nm, 365–405 nm, 480–580 nm, 580–680 nm, 775–900 nm, and 900–1050 nm, respectively, and the corresponding central wavelengths are 345, 385, 529, 628, 835, and 944 nm (Chen et al., 2024b; Yang et al., 2024; Du et al., 2025). Multi-band photometric observations of SN 2024aecx were started from 2024 December 16 (18:45:01 UT). Considering a fast evolution in all bands, we acquired images of the SN field on the consecutive 3 nights in an interval of 1 to 2 hours. A color composite image of the host, indicating the SN location, is shown in Fig. 1.
Currently, Mephisto is in the commissioning phase, equipped with two Andor Technology single-chip (e2v CCD231-C6 6144 6160 sensor) CCD cameras for the blue () and yellow () channels. The red channel () is equipped with a single-chip (e2v CCD 290–99 9216 9232, 10 m sensor) camera developed at the NAOC. The spatial sampling of the red-channel images, 0.286 pixel-1, is better than that of the yellow- and blue-channel images, 0.429 pixel-1. The fields of view of these channels are 43 43. Those will be increased to 2 degrees once the mosaic cameras currently under development are implemented, which is expected to happen in the spring of 2025.
@@
Along with 1.6 m Mephisto, the SN field was also followed up with 50 cm array facilities (two 50 cm telescopes, 50-A and 50-B) located near the Mephisto telescope site (see Chen et al., 2024b). The 50 cm array facilities are equipped with an FLI ML 50100 CCD camera of 8176 6132 pixels (size 6 m). The filter system is the same in all facilities. The 50-A and 50-B telescopes are equipped with and filters and can simultaneously perform observations in and bands, respectively.
@@
The pre-processing of raw frames was performed using a specialized pipeline developed for Mephisto data, which included bias and dark subtraction, flat-fielding, and cosmic ray removal. Afterward, astrometric calibration and photometric measurements were done to extract source information. We used recalibrated Gaia BP/RP (XP) spectra for photometric calibration with the Synthetic Photometry Method (Huang et al., 2024; Xiao et al., 2023). Synthetic AB magnitudes were derived by convolving XP spectra with Mephisto’s transmission curves, the photometric zero-points were determined using high-quality, non-variable stars, correcting for potential gain inaccuracies per CCD output. The calibration uncertainties were better than 0.03 mag in , 0.01 mag in , and 0.005 mag in bands. Since SN 2024aecx is deeply embedded inside its host galaxy, we employed image subtraction to remove the influence of the galaxy background on photometry and construct multiband light curves. The specific steps are as follows:
Date | J.D. | Phase∗ | Grism | Spectral range | Telescope/Instrument |
(yyyy-mm-dd) | (2460000+) | (d) | (Å) | ||
2024–12–17 | 662.37 | 1.92 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2024–12–17 | 662.42 | 1.97 | Gr3 | 3600–9160 | LJT/YFOSC |
2024–12–20 | 665.42 | 4.97 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2024–12–21 | 666.45 | 6.00 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2024–12–24 | 669.35 | 8.90 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2024–12–26 | 671.36 | 10.91 | Gr3 | 3600–9160 | LJT/YFOSC |
2024–12–29 | 674.48 | 14.03 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2024–12–31 | 676.32 | 15.87 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2025–01–06 | 682.47 | 22.02 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2025–01–08 | 684.46 | 23.96 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2025–01–14 | 690.29 | 29.84 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2025–01–15 | 691.30 | 30.85 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
2025–02–07 | 714.41 | 53.96 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
∗ With reference to the date of explosion (JD 2460660.45).
@@
Image alignment: As mentioned previously, the NGC 3521 field was observed multiple times with Mephisto before the SN explosion. For Mephisto, we selected good-quality images with smaller Full Width at Half Maximum (FWHM) and low sky background as reference images in each band. The image frames obtained after the SN explosion were considered target images. For the blue (-band) and yellow channel -band images, we selected 3000 3000 pixel2 image size centered around SN 2024aecx. Similarly, the 4500 4500 pixel2 image portion was selected for the red channel images. For the 50 cm telescope, no pre-explosion images were available. Therefore, we selected a set of high-quality images taken on 27 March 2024, when the transient was too faint to be detected, as reference images. All 50cm images were cropped to 3000 3000 pixel2. Then, science images were aligned relative to the reference images using SWARP (Bertin, 2010).
@@
Image subtraction: The image subtraction was performed with the HOTPANTS software (Becker, 2015). For the differencing, we use a convolution kernel with a half-width set to 2 times the image FWHM for the blue ( band) and yellow ( band) cameras of Mephisto, and the half-width of the substamp extracted around each centroid is set to 5 times the FWHM. For the 50cm telescope, we adopted the same kernel and substamp scaling (2 and 5 times the FWHM, respectively). For the red channel camera of Mephisto, the kernel half-width is set to 3 times the image FWHM, and the substamp extraction was set to 7 times the FWHM. We set the normalization to the reference image and convolve the reference image to match the target image.
Photometry: Aperture photometry was performed on the difference images using SExtractor (Bertin & Arnouts, 1996a) to construct the light curves. Photometric calibration was carried out using the zero-point values derived from the reference image.
Additionally, we retrieved photometric data ( –bands) from the Asteroid Terrestrial-impact Last Alert System (ATLAS Tonry et al., 2018; Smith et al., 2020). It is a large area survey facility with four telescopes situated at different locations. It can perform nightly observation (1 day cadence) of 55 deg2 sky in cyan () and orange () bands, which covers 420–650 nm and 560–820 nm wavelength range. For SN 2024aecx, we downloaded the data from the ATLAS forced photometry service222https://fallingstar-data.com/forcedphot/.
SN 2024aecx was also monitored with Swift/UVOT (Gehrels et al., 2004; Roming et al., 2005) in and bands. These observations start from 2024-12-16 and continue up to 2025-01-13 with good coverage. We obtained the publicly available data from the Swift Archives333https://www.swift.ac.uk/swift_portal/ under the ObsID 00018989. As the SN was deep inside the host, it required the exclusion of the contribution from the host galaxy. Fortunately, Swift had previously monitored the same field under ObsID 00084364. We downloaded this data and used it for the host templates. We utilized Swift_host_subtraction444https://github.com/gterreran/Swift_host_subtraction code (Brown et al., 2009, 2014) to remove the host contribution and perform aperture photometry on the Swift/UVOT images.
@@
∗ Time from the explosion epoch (MJD 60659.95). The maximum possible uncertainty in these estimates is 0.6 days.

### v2

@@ Abstract
We present the results of low-resolution spectroscopic and densely sampled multi-band photometric follow-up of supernova (SN) 2024aecx. The SN was discovered in the spiral galaxy NGC 3521 (distance 11 Mpc) within a day after the explosion. The early spectra of SN 2024aecx show a weak signature of hydrogen lines, which disappeared in 30 days after the explosion. Light curves in all bands show a distinct feature of two peaks, and the first peak is likely due to the shock cooling emission. The early phase light curve evolution of SN 2024aecx has similarity with the typical Type IIb events, but the decay rate in different bands (e.g., = 1.60 0.05 mag, -band) is significantly faster in the post-peak phase. It attained the secondary maximum in 19 days (-band) with a peak absolute magnitude of Mg = –17.94 0.10 mag. SN 2024aecx colors trend redder in early epochs (8 days), followed by a duration in which it grows bluer, then later gets redder again 20 days after explosion. The analytical model fitting to the light curves reveals an envelope mass and progenitor radii in the range of 0.03–0.24 and 169–200 , respectively. Modeling of the pseudo-bolometric light curve suggests that synthesized 56Ni in the explosion was 0.15 M☉ with ejecta mass and kinetic energy of 0.7 M☉ and 0.16 1051 erg, respectively. The observational properties and modeling indicate that the SN 2024aecx progenitor belongs to the extended progenitor category.
@@ 1 Introduction
The evolution of massive stars (10 M⊙) ends when their nuclear fuel is exhausted after the formation of iron group elements in the core (A. Heger et al., 2003; S. E. Woosley & T. A. Weaver, 1986). At this stage, the stellar core collapses (S. Woosley & T. Janka, 2005) due to the strong gravitational force on a very short time scale (of the order of a second), and eventually the star explodes as a core-collapse supernova (CC-SN). Various stellar parameters (e.g., mass, metallicity, mass loss, rotation, binary interaction, etc.) may significantly affect the observational features such as light curves and spectra. Accordingly, the observational properties displayed by these transients at early or maximum light phases make a decisive basis for their classification and subclassification (A. V. Filippenko, 1997). A subgroup (Type IIb, Ib, Ic, and Ic-BL) of these events is termed as ‘stripped-envelope supernovae, SE-SNe’ (R. Minkowski, 1941; A. Clocchiatti & J. C. Wheeler, 1997; A. Gal-Yam, 2017; M. Modjaz et al., 2019).
@@
In SE-SNe, the hydrogen and/or helium envelopes of their progenitor are partially or completely stripped off before the explosion. In particular, hydrogen lines do not appear in SNe Ib, but helium lines are dominant. SNe Ic exhibit neither of these lines in their spectra, and SNe Ic-BL (broad line) are characterized with broad line velocities (15000 km s-1). Type IIb events demonstrate intriguing properties in terms of both spectroscopic and light curve evolution. During the early phases, the hydrogen lines are clearly visible, which disappear progressively in a few weeks (A. V. Filippenko, 1988), and the helium lines dominate the spectra at the later phases (similar to SNe Ib). In addition, double peaks have been observed in the light curves of a handful of IIb SNe, e.g., SN 1993J (M. W. Richmond et al., 1994; R. Barbon et al., 1995), SN 2011dh (I. Arcavi et al., 2011), SN 2011fu (B. Kumar et al., 2013), SN 2013df (A. Morales-Garoffolo et al., 2014; S. D. Van Dyk et al., 2014), SN 2016gkg (L. Tartaglia et al., 2017; M. C. Bersten et al., 2018a), SN 2020bio (C. Pellegrino et al., 2023), SN 2022hnt (J. R. Farah et al., 2025), SN 2024iss (M. Yamanaka et al., 2025), SN 2024uwq (B. M. Subrayan et al., 2025). While the initial peak is mainly interpreted by the cooling of the extended hydrogen-rich envelope in the post-shock breakout phase (see S. W. Falk & W. D. Arnett, 1977a; M. C. Bersten et al., 2012a; E. Nakar & A. L. Piro, 2014; A. L. Piro, 2015; N. Sapir & E. Waxman, 2017; L. Dessart et al., 2018), the secondary peak is mainly powered due to energy released in radioactive decay of 56Ni 56Co 56Fe (W. D. Arnett, 1980). However, a few more scenarios were also proposed to explain the SNe IIb first peak: double 56Ni distribution (M. Orellana & M. C. Bersten, 2022), Thomson scattering and chemical mixing in the SN ejecta (S. H. Park et al., 2024). The double-peak light curve features of SNe IIb have similarities with some ‘Ca-rich’ events which show strong [Ca II] 7291, 7324 Å emission compared to [O II] 6300, 6364 Å (A. V. Filippenko et al., 2003; K. S. Kawabata et al., 2010; H. B. Perets et al., 2010; M. M. Kasliwal et al., 2012). The shape of the first peak can infer the radius and mass of the progenitor (E. Nakar & A. L. Piro, 2014). R. A. Chevalier & A. M. Soderberg (2010) suggested the extended and compact progenitors of SNe IIb that can produce double-peaked and single-peaked light curves, respectively.
@@
From the theoretical modeling aspect, Type IIb SNe progenitors can have both single and binary evolution channels. The radiation-driven stellar wind plays a crucial role in stripping the outer envelope (P. A. Crowther, 2007; C. Georgy, 2012; S.-C. Yoon, 2015) in single massive star models, and mass transfer via a binary companion is a preferred path in the binary stellar evolution models (P. Podsiadlowski, 1992; K. I. Nomoto et al., 1995; J. S. W. Claeys et al., 2011; N. Smith, 2014; N. Sravan et al., 2020; L. Dessart et al., 2024). However, the observational manifestation supports the binary progenitor, which is evident in several direct imaging (e.g., J. R. Maund et al., 2004; G. Folatelli et al., 2014; S. D. Ryder et al., 2018; S. D. Van Dyk et al., 2014, and references therein).
@@
The prevalence of wide-area surveys in recent years has enabled large, sample-based, qualitative studies of transients, particularly SE-SNe (F. Taddia et al., 2015a; Y.-Q. Liu et al., 2016; J. D. Lyman et al., 2016; S. J. Prentice et al., 2016; F. Taddia et al., 2018a; S. J. Prentice et al., 2019; I. Shivvers et al., 2019; M. Williamson et al., 2019; B. Ayala et al., 2025). Better cadence and sampling are equally useful to corroborate whether there is a transition between some SE-SNe, which were classically divided into different types, and recent studies point toward a continuum between them. Especially in a few cases, SNe Type IIb and Ib indicate a thin boundary (e.g., D. Milisavljevic et al., 2013; A. Gangopadhyay et al., 2023; Y. Dong et al., 2024a; N. Yesmin et al., 2025). Therefore, for a better understanding of the observational properties and the progenitor, it is important to study more events in detail, having both common and/or peculiar features.
@@
NGC 3521 is a late-type spiral galaxy SAB(rs)bc with a mixed barred and inner ring morphology. G. de Vaucouleurs et al. (1991, Third Reference Catalogue of Bright Galaxies, RC3), thoroughly examined due to its proximity (11.40 0.56 Mpc)111https://ned.ipac.caltech.edu/. It has an integrated star formation rate of 2.1 M☉ year-1 (A. K. Leroy et al., 2008), and because of the similarity in structural parameters (stellar mass, rotation velocity, optical radius), it is considered as a near structural analogue to the Milky Way (see L. S. Pilyugin et al., 2025, and references therein). The spectroscopic decomposition indicates that NGC 3521 is dominated by the intermediate stellar population (L. Coccato et al., 2018).
SN 2024aecx (ATLAS24rkq) was discovered on MJD 60660.56 (UT 2024 December 16.56) by the ATLAS group in the galaxy NGC 3521 at coordinates (J2000): RA = 11h 05m 4955 DEC = – 00° 02′ 0544 (J. Tonry et al., 2024; H. Stevance et al., 2024). It is one of the very early discovered (within a day) SN after the explosion. The discovery magnitude was 14.68 0.01 (orange-ATLAS). The last non-detection was reported on MJD 60659.35 (UT 2024 December 15.35) with a magnitude of 18.56 mag (orange-ATLAS). AT 2024aecx observations (at MJD = 60660.7) from the Lulin Observatory indicated a fading nature and blue colors ( = – 0.6 mag, = – 0.3 mag) of AT 2024aecx (T. W. Chen et al., 2024). Forced photometry with ATLAS (orange filter) confirmed the fast decline (I. Pérez-Fournon et al., 2024) within a few hours after the explosion. We consider the explosion epoch of SN 2024aecx as MJD 60659.953 (2024 December 15.953), the midpoint of the last non-detection, and the discovery date. The considered explosion time matches reasonably with the estimation from the analytical model fitting as discussed in Section 4. On 2024 December 17, the SN was classified as a Type IIb based on spectra taken from the SNIFS and GMOS instruments mounted on the 88-inch University of Hawaii telescope (J. Hinkle, 2024) and Gemini-North telescope (J. Andrews et al., 2024a). However, Gemini-N/GMOS spectroscopy on 2024 December 19 reclassifies it as Type Ic SN (J. Andrews et al., 2024b). Our initial two spectra were acquired on 2024 December 17 (UT) in a gap of 1 hour (see Section 2). These observations are 8 hours later than J. Andrews et al. (2024a) and J. Hinkle (2024). Detailed spectroscopic investigation (Section 3.4) indicates a weak feature of hydrogen around 6500 Å, which disappears 1 month after the explosion. The double-peak multi-band light curves of SN 2024aecx have a close resemblance to Type IIb events (Section 3.1). Based on these observational characteristics, we consider it a Type IIb SN.
@@ 2.1 Photometry
Multi-band photometric observations of SN 2024aecx were started from 2024 December 16 (18:45:01 UT) with 1.6 m Multi-channel Photometric Survey Telescope (Mephisto, X. Yuan et al., 2020). The facility is located at the Lijiang Observatory and can perform simultaneous observations of a field in three optical bands at a time ( or ). More details about the Mephisto filters and the instrument can be found in recent studies (X. Chen et al., 2024; Y.-P. Yang et al., 2024; Y. Cheng et al., 2025; G. Du et al., 2025; Y.-H. I. Yin et al., 2025, and references therein). Considering a fast evolution in all bands, we acquired images of the SN field on consecutive 3 nights in an interval of 1 to 2 hours. A color composite image of the host, indicating the SN location, is shown in Fig. 1.
@@
Along with 1.6 m Mephisto, the SN field was also followed up with 50 cm array facilities (two 50 cm telescopes, 50-A and 50-B) located near the Mephisto telescope site (see X. Chen et al., 2024). The 50 cm array facilities are equipped with an FLI ML 50100 CCD camera of 8176 6132 pixels (size 6 m). The filter system is the same in all facilities. The 50-A and 50-B telescopes are equipped with and filters, respectively, and can simultaneously perform observations in the and bands.
| MJD | Phase† (days) | Mag | ∗∗ | Filter | Telescope |
| 60660.557 | 0.60 | 14.683 | 0.012 | o | ATLAS |
| 60660.559 | 0.61 | 14.761 | 0.011 | o | ATLAS |
| 60660.563 | 0.61 | 14.700 | 0.009 | o | ATLAS |
| 60660.573 | 0.62 | 14.706 | 0.009 | o | ATLAS |
| 60660.769 | 0.82 | 15.175 | 0.035 | z | 50cm |
| 60660.769 | 0.82 | 15.004 | 0.016 | i | 50cm |
| 60660.772 | 0.82 | 15.225 | 0.081 | z | 50cm |
| 60660.777 | 0.82 | 15.214 | 0.022 | i | 50cm |
| 60660.785 | 0.83 | 15.049 | 0.004 | i | Mephisto |
| 60660.785 | 0.83 | 14.666 | 0.002 | g | Mephisto |
| 60660.788 | 0.83 | 14.694 | 0.002 | g | Mephisto |
| 60660.791 | 0.84 | 15.095 | 0.006 | z | Mephisto |
| 60660.792 | 0.84 | 14.914 | 0.002 | v | Mephisto |
| 60660.794 | 0.84 | 15.083 | 0.006 | z | Mephisto |
| 60660.794 | 0.84 | 14.797 | 0.002 | r | Mephisto |
| 60660.795 | 0.84 | 14.825 | 0.002 | v | Mephisto |
| 60660.798 | 0.85 | 15.044 | 0.007 | z | Mephisto |
| 60660.798 | 0.85 | 14.812 | 0.002 | r | Mephisto |
| 60660.798 | 0.85 | 14.917 | 0.002 | v | Mephisto |
| 60660.857 | 0.90 | 15.532 | 0.115 | UVW1 | Swift |
| 60660.858 | 0.90 | 14.949 | 0.108 | U | Swift |
| 60660.859 | 0.91 | 15.122 | 0.102 | B | Swift |
| 60660.861 | 0.91 | 16.407 | 0.149 | UVW2 | Swift |
| 60660.864 | 0.91 | 15.113 | 0.115 | V | Swift |
∗ This is only a part of the photometric table, and the full table is available as data behind the figure.
† The phase is relative to explosion epoch (MJD 60659.953).
∗∗ Mephisto photometric uncertainties primarily comprise photon noise and detector noise.
@@
The pre-processing of raw frames was performed using a specialized pipeline developed for Mephisto data, which included bias and dark subtraction, flat-fielding, and cosmic ray removal. Afterward, astrometric calibration and photometric measurements were done to extract source information. We used recalibrated Gaia BP/RP (XP) spectra for photometric calibration with the Synthetic Photometry Method (B. Huang et al., 2024; K. Xiao et al., 2023). Synthetic AB magnitudes were derived by convolving XP spectra with Mephisto’s transmission curves; the photometric zero-points were determined using high-quality, non-variable stars, correcting for potential gain inaccuracies per CCD output. The calibration uncertainties were better than 0.03 mag in , 0.01 mag in , and 0.005 mag in bands. Since SN 2024aecx is deeply embedded inside its host galaxy, we employed image subtraction to remove the influence of the galaxy background on photometry and construct multiband light curves. The specific steps are as follows:
@@
Image alignment: As mentioned previously, the NGC 3521 field was observed multiple times with Mephisto before the SN explosion. For Mephisto, we selected good-quality images with smaller Full Width at Half Maximum (FWHM) and low sky background as reference images in each band. The image frames obtained after the SN explosion were considered target images. For the blue (-bands) and yellow channel (-bands) images, we extracted ( 21.5′ 21.5′) cutouts around SN 2024aecx. For the red-channel () images, larger ( 30′ 30′) cutouts were used. For the 50 cm array telescopes, no pre-explosion images were available, so we adopted high-quality images from 2024 March 27 (UT), when the transient was undetectable, as reference frames. All 50 cm images were cropped to the same ( 21.5′ 21.5′) region. Then, science images were aligned relative to the reference images using SWARP (E. Bertin, 2010).
Image subtraction: To examine our data pipeline, implementation of image subtraction and calibration, etc., we initiated a survey of the nearby galaxies, star clusters, and several pilot fields during the commissioning phase. Consequently, the field of NGC 3521 was observed in multi-bands at various epochs since 2023 January 17. Interestingly, the latest Mephisto observations of NGC 3521 were made on UT 2024 December 09, i.e., about 7 days before the SN discovery (see Section 3.1).
@@
The image subtraction was performed with the HOTPANTS software (A. Becker, 2015). For the differencing, we use a convolution kernel with a half-width set to 2 times the image FWHM for the blue ( band) and yellow ( band) cameras of Mephisto, and the half-width of the substamp extracted around each centroid is set to 5 times the FWHM. For the 50cm telescope, we adopted the same kernel and substamp scaling (2 and 5 times the FWHM, respectively). For the red channel camera of Mephisto, the kernel half-width is set to 3 times the image FWHM, and the substamp extraction was set to 7 times the FWHM. We set the normalization to the reference image and convolve the reference image to match the target image.
| MJD | Phase∗ (days) | Grism | Spectral range (Å) | Telescope/Instrument |
| 60661.874 | 1.92 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60661.924 | 1.97 | Gr3 | 3600–9160 | LJT/YFOSC |
| 60664.924 | 4.97 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60665.955 | 6.00 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60668.854 | 8.90 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60670.836 | 10.88 | Gr3 | 3600–9160 | LJT/YFOSC |
| 60673.976 | 14.02 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60675.818 | 15.87 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60681.973 | 22.02 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60683.960 | 24.01 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60688.972 | 29.02 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60689.788 | 29.84 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60690.796 | 30.84 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
| 60713.909 | 53.96 | Gr7, Gr8 | 3500–7800, 5200–9250 | HCT/HFOSC |
∗ With reference to the date of explosion (MJD 60659.953).
@@
∗ Time from the explosion epoch (MJD 60659.953). The maximum possible uncertainty in these estimates is 0.6 days.